### v1

@@
Carlo Alessi & Egidio Falotico, Brain-Inspired Robotics Laboratory, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy.
@@ Abstract
Continuum and soft robots can positively impact diverse sectors, from biomedical applications to marine and space exploration, thanks to their potential to adaptively interact with unstructured environments. However, the complex mechanics exhibited by these robots pose diverse challenges in modeling and control. Reduced order continuum mechanical models based on rod theories have emerged as a promising framework, striking a balance between accurately capturing deformations of slender bodies and computational efficiency. This review paper explores rod-based models and control strategies for continuum and soft robots. In particular, it summarizes the mathematical background underlying the four main rod theories applied in soft robotics. Then, it categorizes the literature on rod models applied to continuum and soft robots based on deformation classes, actuation technology, or robot type. Finally, it reviews recent model-based and learning-based control strategies leveraging rod models. The comprehensive review includes a critical discussion of the trends, advantages, limits, and possible future developments of rod models. This paper could guide researchers intending to simulate and control new soft robots and provide feedback to the design and manufacturing community.
@@
Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in infinite degrees of freedom (DoFs). This characteristic gives rise to a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (Trivedi et al., 2008c). Soft robots, built with soft materials or deformable structures (Laschi et al., 2016), are continuum robots. Generally speaking, soft robots assume a variety of morphologies. In this paper, we refer to both classes with their partial overlap.
@@
Continuum and soft robots deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (Laschi and Cianchetti, 2014; Laschi et al., 2016). Novel design and manufacturing technologies are accelerating the prototyping of highly dexterous robotic systems (Rus and Tolley, 2015; Wang et al., 2024). However, the modeling and control of continuum and soft robots are still challenging problems. Indeed, soft materials cause extreme hyper-redundancy and exhibit nonlinear properties such as hysteresis and stress softening. These challenges inspired the development of diverse modeling techniques (Armanini et al., 2023; Gilbert, 2021), each characterized by underlying assumptions, mathematical framework, and a trade-off between accuracy and computational cost. Soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Fig. 1). Concurrently, innovative control strategies emerged to leverage these models (George Thuruthel et al., 2018; Della Santina et al., 2023).
Popular modeling methods are inherently data-driven (Laschi et al., 2023). Despite not requiring geometric and physical expertise, they rely on pseudo-random motion data to train artificial neural networks mapping actuation to task space (Chin et al., 2020). This robot-independent method effectively derived efficient forward models employed in learned policies for tracking (Centurelli et al., 2022; Piqué et al., 2022) and throwing (Bianchi et al., 2023). However, they are black-box, could suffer overfitting, and the burden of collecting interaction datasets limits their applicability.
Other approaches, like discrete material models, discretize continuous bodies a priori. For example, pseudo-rigid models represent them with a chain of rigid links connected by joints. While these provide satisfactory results for hyper-redundant arms (Venkiteswaran et al., 2019), they poorly approximate elastic structures. In addition, lumped-mass models employ an array of masses, springs, and dampers (Habibi et al., 2020). This modularity can model complex phenomena but could require intensive system identification.
Alternatively, geometrical models utilize geometrical assumptions on the deformed shape of the soft body. For instance, functional models describe the deformations using polynomials (Della Santina and Rus, 2019). In addition, piecewise constant curvature (PCC) models discretize continuous bodies in circular arcs with constant curvature (Xie et al., 2022). These models were effectively used within proportional-derivative (PD) task space controllers for various dynamic tasks (Fischer et al., 2022). Since most continuum robots have actuators parallel to the centerline producing constant-curvature deformations, PCC is viable for uniform and lightweight robots (Webster III and Jones, 2010). However, the suitability of these approximations could degrade when robots are subject to significant external forces or interactions with the environment.
Finally, continuum mechanical models characterize soft robots with continuous configuration spaces and define deformations in physical terms. They enable the simulation of physical interactions and the study of robot mechanics. The classical finite element method (FEM) accurately represents complex 3D geometries to the extent of high computational costs and involved mathematical formulations (Coevoet et al., 2017). Since such models would impede real-time control applications, their use is often limited to designing and simulating soft robotic components (Xavier et al., 2021) or as a benchmark. Nonetheless, recent model order reduction (MOR) techniques are making 3D mechanical models more affordable (Allard et al., 2007; Dubied et al., 2022). Since continuum and soft robots are slender, other methods employ reduced order models of elastic rods, one-dimensional (1D) objects with a dimension (the length) much larger than the others. Elastic rods effectively describe the deformations of slender bodies (i.e., stretching, shearing, bending, twisting), balancing computational efficiency and an accurate representation of complex mechanics. Therefore, rod theories are promising candidates for soft robotics modeling and control, which are the focus of this paper.
1.1 Article Contribution
In this study, we investigate the role of rod theories in the modeling, simulation, and control of continuum and soft robots. The main contributions are:
1.
review of the fundamentals of four major rod theories with a common formulation, which facilitates the comparison and highlights connections to other modeling strategies;
2.
review and analysis of rod models applied to continuum and soft robots, originally classified on deformation classes; and
3.
review and analysis of recent model-based and learning-based controllers leveraging rod models.
First, we summarize the mathematical fundamentals underlying the four main rod theories applied in soft robotics: Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli (Sec. 2). The same notation facilitates finding similarities and differences between the rod theories and supports the analysis of deformation classes. Further, it highlights theoretical insights on spatial discretization (i.e., positions vs configurations), connections to other models (e.g., the polynomial curvature), and implications on the control problem. Also, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec. 3). Then, we review and discuss rod models applied to continuum and soft robots classified by deformation, and eventually robot types or actuation systems (Sec. 4). The analyzed models could guide researchers intending to simulate new soft robots in selecting the most suited rod theory. In addition, it could provide feedback to the design and manufacturing community. Finally, we review and discuss recent model-based and learning-based controllers that leverage rod models (Sec. 5). Along the sections, we critically discuss trends, research gaps, advantages, and limitations of rod models. The paper concludes with an overview of the main remarks (Sec. 6).
1.2 Related Surveys
@@
The seminal paper by Webster III and Jones (2010) reviews PCC kinematic models of continuum robots, emphasizing the robot-specific and robot-independent mappings and providing a perspective on further developments. Conversely, our work focuses on rod models, which generalize PCC approaches in case of significant external forces.
@@
Recently, Armanini et al. (2023) presented a structured overview of all modeling approaches in soft robotics. They propose a classification in four model classes (continuum mechanics, geometrical, discrete material, and surrogate), providing the fundamental theoretical grounds and analyzing their applicability. However, the control problem is only touched marginally. The perspective by Mengaldo et al. (2022) provide a unified formulation for modeling the embodied intelligence of soft robots. It broadly overviews the theories explaining the practical significance of embodied intelligence. Conversely, this paper is a review/survey article specific to rod-based models and controllers. Another work reviews state-of-the-art methods for soft robot modeling and control, emphasizing the use of real-time FEM (Schegg and Duriez, 2022). Also, Zhang et al. (2022) broadly covers a long soft robotics pipeline: design, actuation methods, modeling, and control. It focuses on major technical concerns, providing a perspective on the development of continuum robots. Finally, Sadati et al. (2023) survey MOR techniques and ROM, clarifying their disentangled usage in soft robotics. Within this framework, they broadly overview modeling approaches, constitutive laws, and solution strategies. While they survey the theoretical aspects of modeling, we also provide insights into the experimental validation of rod models.
@@
Regarding control, George Thuruthel et al. (2018) provide a broad overview of model-based and model-free control strategies for continuum soft manipulators, without concern for the modeling. Also, Wang and Chortos (2022) provide a broad review of actuation mechanisms and control strategies, including open-loop, closed-loop, and autonomous methods. They discuss emergent directions in the control-actuator interface, underactuation, and utilization of artificial intelligence. Della Santina et al. (2023) describe the control problem of CSM specifically for the model-based view. They present a unified formulation of soft robot dynamics independent of the modeling technique. Then, they discuss the shape control and tracking problems and explore open challenges (e.g., under-actuation, environment interactions, actuator dynamics, task space control) while surveying the literature. Conversely, we review and discuss rod-based controllers utilizing model-based strategies and deep Reinforcement Learning (RL).
In summary, this review on rod models complements related works with minor intersections.
@@ 2 Background on Rod Theories
An elastic rod is a quasi 1D body in which the length is much larger than the radius of its cross-section. Therefore, the only deformation type considered is along one axis, neglecting other deformations like cross-section deformation. In nature, there are many examples of slender elastic bodies (e.g., hairs, muscle fibers, DNA strands, flagella). In the artificial world, notable examples are continuum and soft robots. In most cases, characterizing the behavior of these robots as that of elastic rods provides a good approximation. The four main Rod Theories used in soft robotics are: Cosserat-Reissner, Kirchoff-Clebsch-Love, Euler-Bernoulli, Timoshenko-Ehrenfest. These models differ in the kinematic assumptions and strain modes considered. An elastic rod might exhibit (i) bending, (ii) torsion, (iii) stretching (i.e., elongation or compression), and (iv) shear. Fig. 2 shows the four rod theories and the modeled strain modes. In this work, we adopted frequently the Lie Algebra notation. The reader can find in Appendix B the definition of the employed operators.

### v2

@@
- GVS
- Geometrical Variable Strain
- ISP
- Implicit Strain Parameterization
@@
- POD
- Proper Orthogonal Decomposition
- SVD
- Singular Value Decomposition
@@
Carlo Alessi, Istituto Italiano di Tecnologia, Genoa, Italy (carlo.alessi@iit.it).
Egidio Falotico, The BioRobotics Institute, Scuola Superiore Sant’Anna, Pisa, Italy (egidio.falotico@santannapisa.it).
@@ Abstract
Continuum and soft robots can transform diverse sectors, including healthcare, agriculture, marine, and space, thanks to their potential to adaptively interact with unstructured environments. These robots exhibit complex mechanics that pose diverse challenges in modeling and control. Among various models, continuum mechanical models based on rod theories can effectively capture the deformations of slender bodies in contact-rich scenarios. This structured review paper focuses on the role of rod models in continuum and soft robot control with a vertical approach. We provide a comprehensive summary of the mathematical background underlying the four main rod theories applied in soft robotics and their variants. Then, we review the literature on rod models applied to continuum and soft robots, providing a novel categorization in deformation classes. Finally, we survey recent model-based and learning-based control strategies leveraging rod models, highlighting their potential in real-world manipulation. We critically discuss the trends, advantages, limitations, research gaps, and possible future developments of rod models. This paper aims to guide researchers who intend to simulate and control new soft robots while providing feedback to the design and manufacturing community.
@@
Continuum robots are defined as robots with distributed deformations throughout their structure, resulting in an infinite number of degrees of freedom (DoFs). This property creates a hyper-redundant configuration space, allowing the robot tip to reach any point in the three-dimensional (3D) workspace with a virtually infinite number of configurations (trivedi2008SoftRob). Soft robots, built with soft materials or deformable structures (laschi2014soft), are a subset of continuum robots. Generally speaking, soft robots can assume a variety of morphologies. In this paper, we refer to both classes, considering their partial overlap.
@@
Continuum and soft robots can deform elastically while safely interacting with unstructured environments thanks to the inherent compliance of soft materials (laschi2016soft). Advances in design and manufacturing technologies are rapidly accelerating the development of highly dexterous robotic systems (rus2015design; WangYanmei2024Advancements). Despite this progress, modeling and controlling these robots remain challenging problems. Indeed, the use of soft materials introduces extreme hyper-redundancy and complex nonlinear behaviors, including hysteresis and stress softening. One of the earliest continuum formulations for hyper-redundant robots represented a manipulator as a smooth backbone curve with distributed parameters and derived partial differential equations for its dynamics (chirikjian1993continuum; chirikjian1994hyper). Since then, diverse modeling techniques were developed (armanini2023soft; gilbert2021mathematical), each characterized by underlying assumptions, mathematical frameworks, and trade-offs between accuracy and computational cost. As summarized in Figure 1, soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical. Concurrently, innovative control strategies emerged to leverage these models (george2018control; della2023model; falotico2024learning). In this paper, we focus on the achievements of rod models.
Rod models were widely employed in continuum and soft robotics, representing a fundamental framework for modeling, simulation, and control. Despite their extensive use, the literature lacks a comprehensive, structured review that consolidates key advancements and critically examines the contributions of these models. To address this gap, this paper investigates the role of rod theories in continuum and soft robotic systems from a vertical perspective. We guide the reader from the mathematical formalization of rod theories, through their application in modeling continuum and soft robots, to recent developments in model-based and learning-based control. By structuring the literature on deformation modes, we aim to provide not only a synthesis of the state of the art but also a practical reference for researchers who want to model continuum and soft robots for control applications.
First, we highlight related surveys in continuum and soft robotics that provide valuable complementary perspectives, where rod models are treated marginally (Sec. 2). Second, we characterize the major modeling techniques to identify their strengths and limitations with respect to rod models (Sec. 3). Then, the main contributions of this paper are as follows:
A comparative review of four major rod theories and discretization techniques, including Cosserat-Reissner, Kirchhoff-Love, Timoshenko, and Euler-Bernoulli. We present them within a consistent mathematical framework to facilitate comparison and highlight connections to alternative modeling approaches (Sec. 4). For completeness, we consider the general modeling of actuation-induced distributed loads with different actuator routings (Sec. 5).
A structured review of rod-based models for continuum and soft robots. We organize rod models into nine deformation classes, providing a solid reference for the development of novel soft robot models (Sec. 6). Deformation modes are the core behaviors that enable soft robots to move and interact with their environment. Classifying the literature by these modes highlights the mechanical effects of actuation, independently of the technology used, and clarifies the modeling and control approaches adopted.
A comprehensive review of model-based control for continuum soft manipulators using rod models. We cover a range of approaches from foundational methods, such as inverse kinematics, inverse dynamics, and feedback linearization, to advanced techniques like model predictive control and optimal control (Sec. 7).
A review of learning-based controllers leveraging rod models. We emphasize Reinforcement Learning (RL) as a promising key enabler to achieve contact-rich manipulation tasks with CSM (Sec. 8).
We provide our insights and critiques along the sections, and finally consolidate the discussion of the main trends, research gaps, and emerging challenges (Sec. 9). While this review focuses on rod theories applied in modeling and control, we believe our insights could inspire new research directions in the wider field of soft robotics. The paper concludes with a summary of the lessons learned (Sec. 10).
2 Related Work
@@
The seminal work of webster2010design reviews kinematic models of piecewise constant curvature (PCC) for continuum robots, highlighting both robot-specific and robot-independent mappings, and outlining directions for future development. In contrast, our work provides a comprehensive treatment of rod models that extend PCC formulations to scenarios involving significant interaction forces, such as during manipulation. Lately, gilbert2021mathematical provides a unified mechanics-based framework for modeling slender biomedical continuum robots, comparing Cosserat rod and general continuum formulations. It offers guidelines for selecting model DoFs to balance accuracy and computational cost, addressing challenges such as under-actuation, actuator–structure coupling, and contact-rich environments. Similarly, dupont2022continuum provides a unified overview on modeling and control of continuum robot architectures—including tendon-actuated, multibackbone, concentric tube, magnetic, and soft designs—focusing on their specific clinical applications while outlining future research directions for widespread clinical use. In parallel, zhang2022survey covers a long soft robotics pipeline, spanning design, actuation methods, modeling, and control. They focus on major technical concerns and provide perspectives on the development of continuum robots. Conversely, the perspective by mengaldo2022concise provides a unified formulation for modeling the embodied intelligence of soft robots. They review the theories explaining the practical significance of embodied intelligence. Conversely, our paper is a vertical investigation of rod theories for the modeling and control of continuum and soft robots. Another work reviews methods for soft robot modeling and control, emphasizing the use of real-time finite element method (FEM) (schegg2022review). Recently, sadati2023reduced focuses on model order reduction (MOR) techniques and reduced order models, clarifying their disentangled usage in soft robotics. Within this framework, they treat modeling approaches, constitutive laws, and solution strategies. While they survey the theoretical aspects of modeling, we also provide practical insights into the experimental validation of rod models. Finally, armanini2023soft presented a structured overview of all modeling approaches in soft robotics. They propose a classification into four model classes (continuum mechanics, geometrical, discrete material, and surrogate), analyzing the fundamental theoretical grounds and their applicability. However, the control problem is only touched marginally.
@@
Regarding control, george2018control provide a broad overview of model-based and model-free control strategies for CSM, however, without concern for the modeling. Building on this, wang2021survey provides a systematic review of Machine Learning (ML) methods for continuum robot control, highlighting model-free approaches for inverse kinematics and closed-loop control, and outlining trends to integrate data-driven methods to handle model uncertainties. Similarly, wang2022control reviews actuation mechanisms and controls, including open-loop, closed-loop, and autonomous methods. They discuss emergent directions in the control-actuator interface, underactuation, and the adoption of artificial intelligence. Subsequently, della2023model describe the problem of model-based control of CSM. They present a unified formulation of soft robot dynamics independent of the modeling technique. Then, they discuss shape control and tracking problems and explore open challenges, such as underactuation, environmental interaction, actuator dynamics, and task-space control, while surveying the literature. Finally, falotico2024learning explore how different models influence the design of learning-based controllers. In contrast to prior work, the present paper walks through the evolution of the literature of controllers for rod models, using both model-based strategies and deep RL.
In summary, our review paper effectively complements related work with minor overlap. Indeed, most existing works mainly provide broad overviews of modeling techniques. In contrast, our vertical focus on rod theories aims to guide the reader from their mathematical formulations to rod-based models of continuum and soft robots, and further to recent control strategies, both model-based and learning-based. We aim to provide new perspectives by structuring the literature around deformation modes. This also serves as a practical reference for modeling continuum robots for control-oriented applications.
3 Features of Soft Robot Models
The soft robotics community proposed various techniques to describe continuum and soft robots. According to armanini2023soft, soft robotics encompasses four model classes: data-driven, discrete, geometrical, and continuum mechanical (Figure 1). Clearly, each strategy has advantages and limitations that depend on the purpose and objectives. Also, some approaches could partially overlap or be combined to derive hybrid models (falotico2024learning). In the context of modeling and control, in this section, we answer the following question:
What are the advantages and disadvantages of rod models with respect to other techniques?
To this end, we provide a concise characterization of the features of the existing modeling techniques and their role in the control problem. In addition, we report the main features in Table 1. In summary, rod models strike a good balance between usability, efficiency, and accuracy.
3.1 Machine Learning-based Models
ML is a powerful data-driven approach to derive the forward model of CSM (laschi2023learning). Unlike analytical methods, which require explicit geometrical and physical descriptions of the robot mechanics, ML only requires training artificial neural networks on descriptive motion data to approximate the map between actuation and task space (chin2020machine). The advantages of this approach include independence from the robot, the computational efficiency of the forward model, and the potential implicit representation of complex physical phenomena, such as hysteresis (falotico2024learning). Indeed, learned forward models were effectively employed within control policies for tracking (centurelli2022closed; pique2022controlling) and throwing (bianchi2023softoss). However, they might become unsuitable for complex manipulation tasks due to the challenges of collecting and labeling representative interaction data (alessi2024pushing). Other limitations include the lack of interpretability of the forward model and the tendency to overfit the training data.
3.2 Geometrical Models
Geometrical models describe the deformed backbone of the CSM with a specific mathematical curve. This idea allows us to define the kinematics of CSM using a finite number of configuration variables that describe the curve and have clear geometrical meaning (e.g., curvature). The most widely used approach, the PCC, discretizes continuous bodies in circular arcs with constant curvature and null torsion (webster2010design). This model is viable for lightweight uniform CSM with actuators parallel to the centerline producing constant curvature. Also, refined methods proposed variable (mahl2014variable) or polynomial curvature (della2019control). Naturally, these methods allow for trade-offs between accuracy and efficiency, and their formulation eases the application of classical rigid-robot controllers. Indeed, geometrical models proved effective for various tracking tasks within model-based (fischer2022dynamic) and learning-based (lou2024controlling) controllers. A limitation of the geometrical assumption, however, is that it can degrade when soft robots are subject to significant external forces from interaction with unstructured environments. Unless augmented with other theories (xie2022simplified), they lack specific parameters to describe materials.
| Features | ML | Geometrical | Discrete | Rod | FEM |
| Computational Efficiency | ✓ | ✓ | ✗ | ||
| Accurate for Complex Geometries | ✗ | ✓ | |||
| Usable in Real-time Control | ✓ | ✓ | ✗ | ||
| Data-driven Adaptability | ✓ | ||||
| Handles Large Deformations | ✓ | ✓ | |||
| Incorporates Material Models | ✗ | ✓ | ✓ | ||
| Ease of Formulation/Implementation | ✓ | ✓ | ✓ | ✗ | |
| Generalization to New Tasks | ✗ | ✓ | ✓ | ||
| Model Interpretability | ✗ | ✓ | ✓ | ✓ | ✓ |
| Handles Physical Interactions | ✓ | ✓ | |||
| (✓) = Highly Suitable or Advantageous | |||||
| () = Moderately Suitable or Partially Effective | |||||
| (✗) = Less Suitable or Challenging | |||||
3.3 Discrete Models
The concept behind discrete models is to discretize the system body at the beginning of the modeling process. Two main methods under this category are the lumped-mass and the pseudo-rigid models.
The lumped-mass method approximates the CSM as a set of lumped masses, springs, and dampers (habibi2020lumped). In this case, the configuration variables represent the spatial displacements of each lumped mass. The modularity of this approach can be adapted to represent arbitrarily complex phenomena (e.g., nonlinear friction) and morphologies (e.g., hybrid kinematic chains). On the other hand, to describe CSM with the same fidelity of continuum mechanics, they require a high number of lumped masses that entail computational effort in addition to data-intensive system identification. Nonetheless, the adoption of this method is facilitated by open-source simulators like SoMo (graule2021somo) and Titan (austin2020titan).
Conversely, the pseudo-rigid approach represents CSM with a chain of rigid links connected by joints, finding an equivalent hyper-redundant rigid robot. Their formulation allows us to exploit the standard controllers from rigid robotics directly (bruno1994robotics). Therefore, pseudo-rigid models provide satisfactory results for hyper-redundant arms (venkiteswaran2019shape). However, they do not perform well in spatial accuracy and would require several DoFs to represent large deformations. Therefore, pseudo-rigid methods are less suitable for modeling the continuous elastic structures found in soft robotics. Nonetheless, morimoto2021model; morimoto2022characterization successfully adapted the MuJoCo simulator (todorov2012mujoco) to describe CSM and train learning-based controllers.
3.4 Continuum Mechanical Models
Finally, continuum mechanical models apply the principles of continuum mechanics to characterize soft robots with continuous configuration spaces and define deformations in physical terms. This formulation enables the simulation of physical interactions (e.g., contact with objects or fluids) and the in-depth study of robot mechanics, particularly for large deformations. These properties make them promising towards CSM interacting with unstructured environments. Depending on the dimensionality and approximation level, we can consider 3D models based on the FEM and low-dimensional formulations such as rod models.
The FEM is one of the most widely used numerical approaches for solving PDE, particularly for modeling 3D elastic bodies. FEM discretizes the soft robot volume into a mesh, which is composed of interconnected finite elements (e.g., line segments, triangles, tetrahedra). The points defining these elements are called nodes, where physical quantities like displacement or stress are computed. Then, the continuous solution can be approximated by interpolating values between the nodes. This approach enables accurate simulation of the complex, nonlinear deformations to the extent of an involved mathematical formulation and high computational costs. Indeed, the solutions provided by standard commercial software (e.g., Abaqus, Ansys, COMSOL) would impede real-time control. Therefore, they were often limited to designing and simulating soft robotic components (xavier2021finite) or as a benchmark. Nonetheless, recent MOR techniques implemented in the SOFA simulator (allard2007sofa; coevoet2017software; dubied2022sim) are making 3D mechanical models more affordable even for learning-based control (agabiti2023whole; menager2023toward).
When the robot is slender, it can be efficiently represented using elastic rod formulations, which approximate the body as a one-dimensional (1D) continuum. In particular, the Cosserat rod theory (CRT) can effectively describe all modes of deformations in CSM (renda2014dynamic; gazzola2018forward; till2019real). The CRT associates a reference system to each cross-section along the CSM. Each cross-section can translate and rotate relative to the adjacent cross-sections to simulate bending, twisting, shearing, and stretching. Thanks to the low-dimensional yet physics-based formulation, rod models retain all the benefits of continuum mechanics at a fraction of the computational cost of FEM. We should note that one limitation of rod models is that they are limited to describing inherently slender structures, although heterogeneous rod assemblies can describe complex systems (renda2018unified; zhang2019modeling). Beyond modeling, rod models are also gaining popularity in model-based control (caradonna2024model) and learning-based manipulation (alessi2024pushing). To sum up, rod theories are promising candidates for soft robotics modeling and control, which are the focus of the remainder of this paper.
@@ 4 Background on Rod Theories
An elastic rod is a quasi 1D body in which the length is much larger than the radius of its cross-section. Therefore, the only deformation type considered is along one axis, neglecting other deformations like cross-section deformation. In nature, there are many examples of slender elastic bodies (e.g., hairs, muscle fibers, DNA strands, flagella). In the artificial world, notable examples are continuum and soft robots. In most cases, characterizing the behavior of these robots as that of elastic rods provides a good approximation. The four main Rod Theories used in soft robotics are: Cosserat-Reissner, Kirchoff-Clebsch-Love, Euler-Bernoulli, Timoshenko-Ehrenfest. These models differ in the kinematic assumptions and strain modes considered. An elastic rod might exhibit (i) bending, (ii) torsion, (iii) stretching (i.e., elongation or compression), and (iv) shear. Figure 2 shows the rod theories and the modeled strains. In this work, we frequently adopted the Lie Algebra notation. Its mathematical operators are defined in Appendix B.